OpenStudy (anonymous):

A mathematics journal has accepted 14 articles for publication. However, due to budgetary restraints only 9 articles can be published this month. How many ways can the journal editor assemble 9 of the articles for publication?

OpenStudy (anonymous):

you are asked to compute the number of ways to choose 9 out of 14 this is written as \(\dbinom{14}{9}\) read "14 choose 9" and computed via \[\frac{14\times 13\times 12\times 11\times 10}{5\times 4\times 3\times 2}\] cancel first, multiply last although this button is probably on your calculator

OpenStudy (anonymous):

ok so what do you mean by cancel first and multiply last

OpenStudy (anonymous):

i mean the answer is a whole number so cancel as must as you can and you will be left only with a multiplication at the end, no division

OpenStudy (anonymous):

for example \(5\times 2=10,3\times 4=12\) so this one becomes \[14\times 13\times 11\]

OpenStudy (anonymous):

I still don't understand? Can you show me step by step

OpenStudy (anonymous):

you can also just cheat http://www.wolframalpha.com/input/?i=14+choose+9

OpenStudy (anonymous):

do you understand how i got \[\frac{14\times 13\times 12\times 11\times 10}{5\times 4\times 3\times 2}\]?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

ok then that is where we have to start

OpenStudy (anonymous):

you are asked how many ways you can choose 9 out of 14 it is easier to answer the question how many ways to choose 5 out of 14 these are obviously the same number, because choosing 9 out of 14 to publish it identical to choosing 5 out of 14 not to publish

OpenStudy (anonymous):

in the numerator i start with 14 and count down until i have 5 numbers that is why the numerator is \[14\times 13\times 12\times 11\times 10\]

OpenStudy (anonymous):

ok im with you

OpenStudy (anonymous):

ohh ok i understand it now thank you so much

OpenStudy (anonymous):

that is like saying 14 choices for the first article 13 choices for the second 12 choices for the third 11 choices for the fourth 10 choices for the fifth

OpenStudy (anonymous):

and if "order counted" you would be done at this step but order does not count in this problem, so i have overcounted by the number of ways i can arrange those 5 articles, which is \(5!=5\times 4\times 3\times 2\)

OpenStudy (anonymous):

that is why the answer is \[\frac{14\times 13\times 12\times 11\times 10}{5\times 4\times 3\times 2}\]

OpenStudy (anonymous):

yep and the final answer is 2,002

OpenStudy (anonymous):

yes it is if you are on a computer you can do what i did in the link above if not, most calculators have a button for this, you can probably find it if you look, or ask someone

OpenStudy (anonymous):

ok thank you so much @satellite73

OpenStudy (anonymous):

yw